sky is clear, forecasts are good for tonight.

17:15 starting darks

19:04 we are open, moving to Kyle's flux std

19:21 collimated. guider avg~0.8". Starting acq.

19:43 starting spec exposure on flux std

19:58 moving to Greg's telluric, guider avg~0.8"

20:31 moving to first science target A773 (greg)

21:20 starting science exposure on A773, we got issues with the mask alignment procedure.

21:46 guider seeing jumped to >2"

22:15 guider seeing avg ~ 2.5"

22:35 moving to A851 (long slit in zJ), seeing very unstable 1.5-2".

23:45 guider in the range 1.5-2.5"

00:10 moving to GOODS-N (Kyle), seeing still bad (>1.5")

02:15 seeing bad, stable(-ish), between 1.5 and 1.7". given seeing, did 1.5" longslit on one object in GOODS-N. now observing Kyle's MOS mask.

03:10 closing (wind > 20 m/sec)

04:48 open, starting acquistion on GOODS-N, seeing ~1.5-1.7"

05:25 observed science field for 10min (guider seeing 1.1"), moving to telluric.

05:45 observing done, will take some flats

-- Main.Clement - 07 Mar 2013
Topic revision: r6 - 07 Mar 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback